OpenStudy (anonymous):

Find the lowest common denominator for the fractions shown? 8/51 & 19/85.....can anyone help me please?

OpenStudy (anonymous):

51 85 / \ / \ 3 17 5 17 LCD is 5*3*17 = ?

OpenStudy (anonymous):

lcm of fraction=(lcm of num)/(hcf of denom) here lcm of num=(19*8)=152 hcf of denom =17 hence reqd ans =152/17

OpenStudy (anonymous):

Thanks :)